### v1

@@ I Introduction
Brain-Computer Interfaces (BCIs) establish direct communication pathways between the central nervous system and external devices [schalk2004bci2000, wu2013convergence, zhang2019human]. Electroencephalography (EEG) serves as the predominant non-invasive modality for BCI systems, capable of capturing macroscopic cortical activity with millisecond-level temporal resolution, high portability, and cost-efficiency. Consequently, EEG has become the cornerstone of critical applications such as motor imagery [altaheri2021deep, dai2020hs], affective state recognition [dadebayev2022eeg, gao2024multimodal, zhao2025wearable], epilepsy diagnosis [ahmad2022eeg, yildiz2022unsupervised], sleep staging [phan2021xsleepnet, wang2024generalizable, zhou2024personalized], consciousness assessment [bai2021managing, zhao2024docter], and depression detection [yasin2021eeg, wang2025m]. However, achieving universal BCI remains challenging, impeded by the multifaceted heterogeneity of neural data. At the signal level, recordings are inherently non-stationary and suffer from low signal-to-noise ratios, stemming from anatomical variability [saha2020intra] and pervasive artifacts like muscle activity and line noise [bjork1953electrical, mannan2018identification, uriguen2015eeg, jamil2021artifact]. Physically, diverse acquisition hardware employing varying channel counts and reference schemes induces inconsistent spatial alignments across datasets [acharya2019overview]. Compounding these discrepancies is the functional heterogeneity of BCI paradigms themselves, which recruit divergent neural mechanisms ranging from frequency-specific sensorimotor rhythms to complex, widespread network oscillations. These factors collectively prevent cross-scenario generalization, necessitating extensive adaptation to novel deployment contexts and hindering the realization of universal BCI systems.
@@
Inspired by the success of large language and vision models [bommasani2021opportunities, min2023recent, awais2025foundation, xu2025lvlm], the field has increasingly turned toward EEG foundation models to learn universal representations from vast unlabeled corpora [zhou2025brain, li2025foundation, gu2025foundation, lai2025simple, yang2025foundation]. These approaches typically employ self-supervised strategies like masked signal reconstruction. Despite their statistical capacity, these data-driven models rely on end-to-end fine-tuning to adapt to downstream tasks yet frequently exhibit limited efficacy under a frozen-probing protocol [lee2025are, xiong2025eeg, wu2025adabrain]. We argue that establishing robust performance in this frozen setting, where the pretrained backbone remains fixed, is a prerequisite for universal BCI. It serves as a critical benchmark for validating whether learned representations capture genuinely invariant neurophysiological mechanisms rather than scenario-specific statistics. We attribute the current deficiency to the reliance on generic sequence modeling assumptions. By prioritizing data-driven patterns while underutilizing the distinct biophysical and dynamical principles governing neural activity, existing architectures yield representations that lack intrinsic universality.
@@
This limitation stems from neglecting the structured physical processes that define EEG. Two core principles shape this structure. First, spatial organization is dictated by volume conduction, where passive current spread creates spatially blurred, distance-dependent correlations among electrodes [nunez2006electric]. Second, temporal organization exhibits a hierarchy spanning from millisecond-level oscillations to second-level cognitive adaptations [buzsaki2006rhythms, fairhall2001efficiency]. Furthermore, from a dynamical systems perspective, macroscopic brain function is governed by low-dimensional order parameters [strogatz2024nonlinear, breakspear2017dynamic]. Key statistics like spectral power [buzsaki2006rhythms], functional connectivity [varela2001brainweb], cross-frequency coupling [canolty2006high], and dynamic complexity [richman2000physiological] constitute the stable, interpretable features of brain states. Neglecting these priors yields neurobiologically agnostic representations, which necessitates extensive parameter updates to recover meaningful features, thereby hindering the emergence of the intrinsic universality required for robust frozen-probing.
@@ II-A Traditional EEG Decoding
Electroencephalography (EEG) decoding aims to infer cognitive, affective, or pathological states from scalp-recorded electrical activity. Early methodologies predominantly relied on handcrafted features, including time-domain statistics, spectral power, and connectivity measures, coupled with classifiers such as support vector machines or linear discriminant analysis [mcfarland2006bci, lotte2007review, wu2014probabilistic, wang2023sparse]. These approaches necessitate extensive domain expertise and often exhibit limited generalization across diverse subjects and recording conditions.
The emergence of deep learning has shifted the field toward end-to-end representation learning [lecun2015deep]. Numerous studies now employ deep architectures to automate feature extraction from raw or minimally processed EEG [parvaneh2019cardiac, craik2019deep, al2021deep, sekkal2022automatic, yangpyhealth]. Convolutional Neural Networks (CNNs) have been widely adopted to capture local spatiotemporal patterns, with architectures such as DeepConvNet [schirrmeister2017deep], C2CM [sakhavi2018c2cm], EEGNet [lawhern2018eegnet], and TSception [ding2022tsception] demonstrating efficacy in motor imagery, emotion recognition, and seizure detection [abdelhameed2021deep]. To model long-range temporal dependencies, Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have been applied to tasks including sleep staging and BCI control [wang2018lstm, phan2019seqsleepnet]. Hybrid CNN-LSTM models further combine local feature extraction with sequential modeling for applications such as sleep staging and emotion recognition [supratak2017deepsleepnet, zhang2019cnnlstm, dar2020cnn, li2022motor, wang20232d].
More recently, the Transformer architecture has been introduced to EEG analysis, utilizing self-attention to model global spatiotemporal interactions without fixed convolutional inductive biases [song2021transformer, liu2021transformers, du2022eeg, phan2022sleeptransformer]. Several works integrate CNNs and Transformers to jointly exploit local invariance and global context [song2022eeg, peh2022transformer, wang2023narcolepsy, zhou2024simplifying, wang2024caresleepnet]. Graph Neural Networks (GNNs) have also been explored to explicitly model inter-channel relationships using predefined or learned connectivity structures [jia2020graphsleepnet, jia2021multi, ding2023lggnet]. Despite success on specific benchmarks, these supervised methods face fundamental limitations: the high cost of acquiring labeled data and poor cross-dataset generalization due to variations in hardware, montage, and physiology. This has motivated the shift toward self-supervised pretraining.
@@ II-B EEG Foundation Models
Foundation models have revolutionized vision, language, and multimodal learning [bommasani2021opportunities, devlin2018bert, he2022masked, radford2021learning, achiam2023gpt, kirillov2023segment, videoworldsimulators2024]. Extending this paradigm to time series, general-purpose models have been proposed for domains including climate, finance, and traffic [eldele2021time, zhang2022self, woo2024unified, chen2024visionts, deng2024lpsgm], with a subset targeting neural signals.
In the EEG domain, early self-supervised efforts focused on pretext tasks to extract invariant representations from unlabeled data [banville2021uncovering]. BENDER [kostas2021bendr] employs contrastive learning on generic EEG embeddings, while MAEEG [chien2022maeeg] utilizes a masked autoencoder framework for signal reconstruction. Subsequent works extended these concepts to intracranial recordings: BrainBERT [wang2022brainbert] reconstructs masked spectrograms of stereo-EEG, and Brant [zhang2023brant] introduces a Transformer-based model for long-term dependency modeling. Its extensions, Brant-2 [yuan2024brant] and Brant-X [zhang2024brant], further incorporate scalp EEG and multi-modal physiological alignment.
Recent research has increasingly targeted scalp EEG. BIOT [yang2023biot] employs a linear transformer for joint pretraining across diverse biosignals. EEG2Rep [foumani2024eeg2rep] predicts masked segments in a latent space, and LaBraM [jiang2024large] introduces neural token prediction for EEG patches. Neuro-GPT [cui2024neuro] utilizes a decoder-only architecture with autoregressive pretraining, while EEGPT [wang2024eegpt] employs a dual objective for universal feature extraction. Approaches such as NeuroLM [jiang2025neurolm], UniMind [lu2025unimind], and ELASTIQ [jiang2025elastiq] explore harnessing Large Language Models (LLMs) by treating EEG as a foreign language. Architecturally, CBraMod [wang2025cbramod] uses criss-cross attention with asymmetric positional encoding, EEGMamba [wang2025eegmamba] and CodeBrain [ma2025codebrain] leverage state space models for spatiotemporal dependencies, and LUNA [doner2025luna] unifies diverse electrode layouts via learned queries. Other innovations include mixture-of-experts in NeurIPT [fang2025neuript], structured sparse attention in CSBrain [zhou2025csbrain], discrete tokenization in BrainOmni [xiao2025brainomni], and massive-scale pretraining in REVE [ouahidi2025reve].
While these models improve transferability, they predominantly adapt generic architectures from vision or language. Lacking explicit biophysical constraints, they rely heavily on end-to-end fine-tuning and typically struggle under frozen-probing protocols [lee2025are, xiong2025eeg, wu2025adabrain].
@@ III-D Neurophysiologically Grounded Positional Encodings
Standard positional encodings in vision or language models assume either fixed geometric priors (e.g., 2D grid in ViT [dosovitskiy2020image]) or purely learnable embeddings. While sufficient for many domains, EEG is governed by distinct biophysical laws. Explicitly modeling these principles can provide stronger inductive biases than generic learnable embeddings. To address this, we introduce two neurophysiologically grounded positional encodings that explicitly embed the physics of EEG generation into the model architecture.
Channel Positional Encoding: Modeling Volume Conduction. Scalp EEG potentials arise from the passive spread of electrical currents through the head’s conductive tissues, a process known as volume conduction [nunez2006electric]. Under the quasi-static approximation of Maxwell’s equations, the potential $\phi(\mathbf{r})$ at scalp location $\mathbf{r}$ due to a current dipole $\mathbf{q}$ at source location $\mathbf{r}_{s}$ is:
@@
To model this, we leverage 3D electrode coordinates $\mathbf{P}_{\text{ch}}\in\mathbb{R}^{C\times 3}$ in a standardized anatomical space (e.g., MNI, in meters). Pairwise Euclidean distances are computed as $\mathbf{D}_{ij}=\|\mathbf{p}_{i}-\mathbf{p}_{j}\|_{2}$. Inspired by the exponential decay observed in volume conduction models [grech2008review], we define a learnable spatial kernel:
@@
From a biophysical perspective, EEG can be modeled as a linear mixture $\mathbf{x}=\mathbf{A}\mathbf{s}+\boldsymbol{\epsilon}$, where $\mathbf{A}\in\mathbb{R}^{C\times Q}$ is the lead field matrix and $\mathbf{s}$ denotes source activity. Under isotropic and uncorrelated sources ($\mathbb{E}[\mathbf{s}\mathbf{s}^{\top}]=\mathbf{I}$), the covariance of observed signals satisfies $\Sigma_{ij}\propto\exp(-\|\mathbf{p}_{i}-\mathbf{p}_{j}\|/\tau)$ [nunez2006electric]. Our kernel $\mathbf{K}_{ij}$ thus approximates this expected spatial correlation structure. To ensure stable aggregation, we normalize each row of the kernel to form a convex combination:
@@
Temporal Positional Encoding: Capturing Slow Modulations and Adaptation.   While temporal convolutional patch embeddings capture fine-grained oscillations, higher-order brain states involve slower modulations spanning seconds to minutes, including alpha power fluctuations and adaptation to sustained stimuli [buzsaki2006rhythms, fairhall2001efficiency]. These dynamics provide critical temporal context. However, generic positional encodings, including standard sinusoidal and learnable embeddings, treat time indices as symmetric dimensions or arbitrary markers. They lack inductive biases for the ”arrow of time,” failing to capture the dissipative nature of neural adaptation. To model this cross-second structure, we design a temporal positional encoding operating at second-level granularity that combines two biologically inspired bases:
@@
This is grounded in the observation that neural responses to sustained input often follow $r(t)\propto e^{-t/\tau}$ [fairhall2001efficiency], where $\tau=1/d_{m}$ is the adaptation time constant. The decay rates are pre-defined to span $\tau\in[1,100]$ seconds, covering sensory adaptation (seconds) to cognitive maintenance (tens of seconds), and are held fixed during training.
@@ III-E Spatiotemporal Transformer Encoder
The embedded spatiotemporal tokens $\mathbf{Z}^{(0)}\in\mathbb{R}^{N\times D}$, with $N=C\times S$ patches and embedding dimension $D$, are processed by a stack of $L$ Transformer encoder layers. While we adopt the standard pre-normalization architecture [wang2019learning], its suitability for EEG stems from a fundamental alignment with the brain’s own computational principles: neural processing is inherently distributed (across regions) and dynamic (across time scales). The self-attention mechanism naturally models such long-range, flexible interactions, mirroring how distant brain areas coordinate via phase synchronization or cross-frequency coupling during cognition.
Formally, each layer follows the pre-normalization scheme [wang2019learning]:

### v2

@@ I Introduction
Brain-Computer Interfaces (BCIs) establish direct communication pathways between the central nervous system and external devices [Schalk et al., 2004; Wu et al., 2013; Zhang et al., 2019b]. Electroencephalography (EEG) serves as the predominant non-invasive modality for BCI systems, capable of capturing macroscopic cortical activity with millisecond-level temporal resolution, high portability, and cost-efficiency. Consequently, EEG has become the cornerstone of critical applications such as motor imagery [Altaheri et al., 2021; Dai et al., 2020], affective state recognition [Dadebayev et al., 2022; Gao et al., 2024; Zhao et al., 2025], epilepsy diagnosis [Ahmad et al., 2022; Yıldız et al., 2022], sleep staging [Phan et al., 2021; Wang et al., 2024b; Zhou et al., 2024a], consciousness assessment [Bai et al., 2021; Zhao et al., 2024], and depression detection [Yasin et al., 2021; Wang et al., 2025c]. However, achieving universal BCI remains challenging, impeded by the multifaceted heterogeneity of neural data. At the signal level, recordings are inherently non-stationary and suffer from low signal-to-noise ratios, stemming from anatomical variability [Saha and Baumert, 2020] and pervasive artifacts like muscle activity and line noise [Björk and Kugelberg, 1953; Mannan et al., 2018; Urigüen and Garcia-Zapirain, 2015; Jamil et al., 2021]. Physically, diverse acquisition hardware employing varying channel counts and reference schemes induces inconsistent spatial alignments across datasets [Acharya and Acharya, 2019]. Compounding these discrepancies is the functional heterogeneity of BCI paradigms themselves, which recruit divergent neural mechanisms ranging from frequency-specific sensorimotor rhythms to complex, widespread network oscillations. These factors collectively prevent cross-scenario generalization, necessitating extensive adaptation to novel deployment contexts and hindering the realization of universal BCI systems.
@@
Inspired by the success of large language and vision models [Bommasani et al., 2021; Min et al., 2023; Awais et al., 2025; Xu et al., 2025], the field has increasingly turned toward EEG foundation models to learn universal representations from vast unlabeled corpora [Zhou et al., 2025a; Li et al., 2025; Gu et al., 2025; Lai et al., 2025; Yang et al., 2025]. These approaches typically employ self-supervised strategies like masked signal reconstruction. Despite their statistical capacity, these data-driven models rely on end-to-end fine-tuning to adapt to downstream tasks yet frequently exhibit limited efficacy under a frozen-probing protocol [Lee et al., 2025; Xiong et al., 2025; Wu et al., 2025]. We argue that establishing robust performance in this frozen setting, where the pretrained backbone remains fixed, is a prerequisite for universal BCI. It serves as a critical benchmark for validating whether learned representations capture genuinely invariant neurophysiological mechanisms rather than scenario-specific statistics. We attribute the current deficiency to the reliance on generic sequence modeling assumptions. By prioritizing data-driven patterns while underutilizing the distinct biophysical and dynamical principles governing neural activity, existing architectures yield representations that lack intrinsic universality.
@@
This limitation stems from neglecting the structured physical processes that define EEG. Two core principles shape this structure. First, spatial organization is dictated by volume conduction, where passive current spread creates spatially blurred, distance-dependent correlations among electrodes [Nunez and Srinivasan, 2006]. Second, temporal organization exhibits a hierarchy spanning from millisecond-level oscillations to second-level cognitive adaptations [Buzsáki, 2006; Fairhall et al., 2001]. Furthermore, from a dynamical systems perspective, macroscopic brain function is governed by low-dimensional order parameters [Strogatz, 2024; Breakspear, 2017]. Key statistics like spectral power [Buzsáki, 2006], functional connectivity [Varela et al., 2001], cross-frequency coupling [Canolty et al., 2006], and dynamic complexity [Richman and Moorman, 2000] constitute the stable, interpretable features of brain states. Neglecting these priors yields neurobiologically agnostic representations, which necessitates extensive parameter updates to recover meaningful features, thereby hindering the emergence of the intrinsic universality required for robust frozen-probing.
@@ II-A Traditional EEG Decoding
Electroencephalography (EEG) decoding aims to infer cognitive, affective, or pathological states from scalp-recorded electrical activity. Early methodologies predominantly relied on handcrafted features, including time-domain statistics, spectral power, and connectivity measures, coupled with classifiers such as support vector machines or linear discriminant analysis [McFarland et al., 2006; Lotte et al., 2007; Wu et al., 2014; Wang et al., 2023c]. These approaches necessitate extensive domain expertise and often exhibit limited generalization across diverse subjects and recording conditions.
The emergence of deep learning has shifted the field toward end-to-end representation learning [LeCun et al., 2015]. Numerous studies now employ deep architectures to automate feature extraction from raw or minimally processed EEG [Parvaneh et al., 2019; Craik et al., 2019; Al-Saegh et al., 2021; Sekkal et al., 2022; Yang et al., 2022]. Convolutional Neural Networks (CNNs) have been widely adopted to capture local spatiotemporal patterns, with architectures such as DeepConvNet [Schirrmeister et al., 2017], C2CM [Sakhavi et al., 2018], EEGNet [Lawhern et al., 2018], and TSception [Ding et al., 2022] demonstrating efficacy in motor imagery, emotion recognition, and seizure detection [Abdelhameed and Bayoumi, 2021]. To model long-range temporal dependencies, Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have been applied to tasks including sleep staging and BCI control [Wang et al., 2018; Phan et al., 2019]. Hybrid CNN-LSTM models further combine local feature extraction with sequential modeling for applications such as sleep staging and emotion recognition [Supratak et al., 2017; Zhang et al., 2019a; Dar et al., 2020; Li et al., 2022; Wang et al., 2023a].
More recently, the Transformer architecture has been introduced to EEG analysis, utilizing self-attention to model global spatiotemporal interactions without fixed convolutional inductive biases [Song et al., 2021; Liu et al., 2021a; Du et al., 2022; Phan et al., 2022]. Several works integrate CNNs and Transformers to jointly exploit local invariance and global context [Song et al., 2022; Peh et al., 2022; Wang et al., 2023b; Zhou et al., 2024b; Wang et al., 2024c]. Graph Neural Networks (GNNs) have also been explored to explicitly model inter-channel relationships using predefined or learned connectivity structures [Jia et al., 2020, 2021; Ding et al., 2023]. Despite success on specific benchmarks, these supervised methods face fundamental limitations: the high cost of acquiring labeled data and poor cross-dataset generalization due to variations in hardware, montage, and physiology. This has motivated the shift toward self-supervised pretraining.
@@ II-B EEG Foundation Models
Foundation models have revolutionized vision, language, and multimodal learning [Bommasani et al., 2021; Devlin et al., 2018; He et al., 2022; Radford et al., 2021; Achiam et al., 2023; Kirillov et al., 2023; Brooks et al., 2024]. Extending this paradigm to time series, general-purpose models have been proposed for domains including climate, finance, and traffic [Eldele et al., 2021; Zhang et al., 2022; Woo et al., 2024; Chen et al., 2024; Deng et al., 2024], with a subset targeting neural signals.
In the EEG domain, early self-supervised efforts focused on pretext tasks to extract invariant representations from unlabeled data [Banville et al., 2021]. BENDER [Kostas et al., 2021] employs contrastive learning on generic EEG embeddings, while MAEEG [Chien et al., 2022] utilizes a masked autoencoder framework for signal reconstruction. Subsequent works extended these concepts to intracranial recordings: BrainBERT [Wang et al., 2022] reconstructs masked spectrograms of stereo-EEG, and Brant [Zhang et al., 2023] introduces a Transformer-based model for long-term dependency modeling. Its extensions, Brant-2 [Yuan et al., 2024] and Brant-X [Zhang et al., 2024a], further incorporate scalp EEG and multi-modal physiological alignment.
Recent research has increasingly targeted scalp EEG. BIOT [Yang et al., 2023] employs a linear transformer for joint pretraining across diverse biosignals. EEG2Rep [Foumani et al., 2024] predicts masked segments in a latent space, and LaBraM [Jiang et al., 2024b] introduces neural token prediction for EEG patches. Neuro-GPT [Cui et al., 2024] utilizes a decoder-only architecture with autoregressive pretraining, while EEGPT [Wang et al., 2024a] employs a dual objective for universal feature extraction. Approaches such as NeuroLM [Jiang et al., 2025b], UniMind [Lu et al., 2025], and ELASTIQ [Jiang et al., 2025a] explore harnessing Large Language Models (LLMs) by treating EEG as a foreign language. Architecturally, CBraMod [Wang et al., 2025a] uses criss-cross attention with asymmetric positional encoding, EEGMamba [Wang et al., 2025b] and CodeBrain [Ma et al., 2025] leverage state space models for spatiotemporal dependencies, and LUNA [Döner et al., 2025] unifies diverse electrode layouts via learned queries. Other innovations include mixture-of-experts in NeurIPT [Fang et al., 2025], structured sparse attention in CSBrain [Zhou et al., 2025b], discrete tokenization in BrainOmni [Xiao et al., 2025], and massive-scale pretraining in REVE [Ouahidi et al., 2025].
While these models improve transferability, they predominantly adapt generic architectures from vision or language. Lacking explicit biophysical constraints, they rely heavily on end-to-end fine-tuning and typically struggle under frozen-probing protocols [Lee et al., 2025; Xiong et al., 2025; Wu et al., 2025].
@@ III-D Neurophysiologically Grounded Positional Encodings
Standard positional encodings in vision or language models assume either fixed geometric priors (e.g., 2D grid in ViT [Dosovitskiy et al., 2020]) or purely learnable embeddings. While sufficient for many domains, EEG is governed by distinct biophysical laws. Explicitly modeling these principles can provide stronger inductive biases than generic learnable embeddings. To address this, we introduce two neurophysiologically grounded positional encodings that explicitly embed the physics of EEG generation into the model architecture.
Channel Positional Encoding: Modeling Volume Conduction. Scalp EEG potentials arise from the passive spread of electrical currents through the head’s conductive tissues, a process known as volume conduction [Nunez and Srinivasan, 2006]. Under the quasi-static approximation of Maxwell’s equations, the potential $\phi(\mathbf{r})$ at scalp location $\mathbf{r}$ due to a current dipole $\mathbf{q}$ at source location $\mathbf{r}_{s}$ is:
@@
To model this, we leverage 3D electrode coordinates $\mathbf{P}_{\text{ch}}\in\mathbb{R}^{C\times 3}$ in a standardized anatomical space (e.g., MNI, in meters). Pairwise Euclidean distances are computed as $\mathbf{D}_{ij}=\|\mathbf{p}_{i}-\mathbf{p}_{j}\|_{2}$. Inspired by the exponential decay observed in volume conduction models [Grech et al., 2008], we define a learnable spatial kernel:
@@
From a biophysical perspective, EEG can be modeled as a linear mixture $\mathbf{x}=\mathbf{A}\mathbf{s}+\boldsymbol{\epsilon}$, where $\mathbf{A}\in\mathbb{R}^{C\times Q}$ is the lead field matrix and $\mathbf{s}$ denotes source activity. Under isotropic and uncorrelated sources ($\mathbb{E}[\mathbf{s}\mathbf{s}^{\top}]=\mathbf{I}$), the covariance of observed signals satisfies $\Sigma_{ij}\propto\exp(-\|\mathbf{p}_{i}-\mathbf{p}_{j}\|/\tau)$ [Nunez and Srinivasan, 2006]. Our kernel $\mathbf{K}_{ij}$ thus approximates this expected spatial correlation structure. To ensure stable aggregation, we normalize each row of the kernel to form a convex combination:
@@
Temporal Positional Encoding: Capturing Slow Modulations and Adaptation.   While temporal convolutional patch embeddings capture fine-grained oscillations, higher-order brain states involve slower modulations spanning seconds to minutes, including alpha power fluctuations and adaptation to sustained stimuli [Buzsáki, 2006; Fairhall et al., 2001]. These dynamics provide critical temporal context. However, generic positional encodings, including standard sinusoidal and learnable embeddings, treat time indices as symmetric dimensions or arbitrary markers. They lack inductive biases for the ”arrow of time,” failing to capture the dissipative nature of neural adaptation. To model this cross-second structure, we design a temporal positional encoding operating at second-level granularity that combines two biologically inspired bases:
@@
This is grounded in the observation that neural responses to sustained input often follow $r(t)\propto e^{-t/\tau}$ [Fairhall et al., 2001], where $\tau=1/d_{m}$ is the adaptation time constant. The decay rates are pre-defined to span $\tau\in[1,100]$ seconds, covering sensory adaptation (seconds) to cognitive maintenance (tens of seconds), and are held fixed during training.
@@ III-E Spatiotemporal Transformer Encoder
The embedded spatiotemporal tokens $\mathbf{Z}^{(0)}\in\mathbb{R}^{N\times D}$, with $N=C\times S$ patches and embedding dimension $D$, are processed by a stack of $L$ Transformer encoder layers. While we adopt the standard pre-normalization architecture [Wang et al., 2019], its suitability for EEG stems from a fundamental alignment with the brain’s own computational principles: neural processing is inherently distributed (across regions) and dynamic (across time scales). The self-attention mechanism naturally models such long-range, flexible interactions, mirroring how distant brain areas coordinate via phase synchronization or cross-frequency coupling during cognition.
Formally, each layer follows the pre-normalization scheme [Wang et al., 2019]: